# DenseVLM: A Retrieval and Decoupled Alignment Framework for Open-Vocabulary Dense Prediction

### v1

@@ 4.2 Ablation study
4.3 Comparison with other VLMs
Quantitative evaluation. We perform a comprehensive quantitative evaluation of dense representations across multiple VLMs on the COCO Panoptic [29] and ADE20K Panoptic [58] datasets. Specifically, we compare methods that are trained on large-scale image-text pairs, including OpenCLIP [3], DFN [10], SigLIP [55], and EVA-CLIP [41]. As shown in Tab. 5, while these methods achieve strong zero-shot image classification performance, their performance in region recognition is notably suboptimal. For instance, EVA-CLIP achieves only a Top-1 box classification accuracy of 44.3% on COCO and 33.0% on ADE20K. Although RegionCLIP [57], which is trained on region-text pairs, shows improved performance on COCO, but fails to generalize to datasets such as ADE20K. Moreover, CLIPSelf [46], which incorporates self-distillation, achieves higher box classification accuracy but performs poorly in mask classification. In contrast, our method, DenseVLM, notably outperforms these models, achieving a 4.3% improvement in Top-1 Mask-T accuracy and a 3.9% improvement in Top-1 Mask-S accuracy on COCO, highlighting its effectiveness in dense prediction tasks.
Qualitative results. We also visualize the attention maps generated by various methods in response to text descriptions of object categories by employing cosine similarity. As illustrated in Fig. 4, our DenseVLM exhibits superior accuracy and completeness in object localization compared to EVA-CLIP and CLIPSelf. Furthermore, DenseVLM is less affected by interference from other objects, effectively maintaining semantic separation, even when their visual features are fused within the network.
@@ 4.4 Zero-shot comparisons at different data scales
ViT-B/16 CLIPSelf 67.6 64.4 44.5
ViT-B/16 DenseVLM 71.9 70.0 47.8
ViT-B/16* EVA-CLIP 44.3 44.7 26.2
ViT-B/16* CLIPSelf 69.1 66.7 41.7
ViT-B/16* DenseVLM 73.4 71.0 45.6
@@
Table 6: Results and comparisons of different backbones. * indicates the model initialized by EVA-CLIP [41].
Method Backbone A-150 A-847 PC-459
PACL [38] ViT-B/16 31.4 - -
OVSeg [27] ViT-B/16 24.8 7.1 11.0
MAFT [17] ConvNeXt-B 29.1 10.1 12.6
SED [48] ConvNeXt-B 31.6 11.4 18.6
SCAN [31] ViT-B/16 30.8 10.8 13.2
SAN [50] ViT-B/16 27.4 10.0 13.0
SAN+DenseVLM ViT-B/16 29.5+2.1 10.4+0.4 15.6+2.6
CAT-Seg [4] ViT-B/16 31.4 11.7 18.4
CAT-Seg+DenseVLM ViT-B/16 34.1+2.1 12.1+0.4 18.7+0.3
@@
4.5 Extending to different backbones
4.6 Application to open-vocabulary dense tasks
@@ 5 Conclusions
  • In Sec. C, we offer supplementary experiments on input image sizes, region proposals, backbones and category sets for training VLMs.
@@ Appendix A Limitations and broader impact
Limitations: Our aim is to develop a region-language alignment model that effectively integrates local visual and semantic features, thereby improving open-vocabulary dense prediction performance. Compared to previous pre-trained Vision-Language Models (VLMs) [3, 41, 57, 46], our proposed DenseVLM achieves superior results and significantly improves downstream task performance. We believe DenseVLM has even greater potential. 1) Scalability. DenseVLM is designed within an efficient, unsupervised region-language alignment framework, making it adaptable to various datasets. However, computational resource limitations have restricted our ability to scale to larger datasets. 2) Model capacity. We employ the ViT-L/14 model from CLIPSelf [46] as a powerful Pre-trained VLM (P-VLM). Utilizing more robust VLMs can yield better performance, and transferring their rich semantic knowledge to lightweight models is a promising direction. 3) Fine-grained Semantic. We categorize objects into broad thing and stuff classes. Fine-grained semantic segmentation and decoupled alignment would enhance the model’s ability to distinguish between similar categories. We plan to explore these avenues in our future research.
@@ Appendix C Additional experiments
Ablation study on input image sizes. To evaluate the effect of input image size on DenseVLM, we conduct experiments with five distinct resolutions: 224, 320, 512, 768 and 1024 pixels, for both training and inference. As shown in Tab. 10, the model’s performance on the region classification task improves as the image resolution increases from 224 to 1024 pixels. This enhancement can be attributed to the greater detail captured at higher resolutions. However, this improvement comes at the cost of significantly increased GPU memory usage. However, this improvement comes at the cost of significantly increased GPU memory usage. Considering the trade-off between computational resources and model performance, we resize the images to 512$\times$512 pixels to achieve an optimal balance.
Ablation study on using region proposals. Following RegionCLIP [57] for fine-tuning VLMs with pseudo-labelled region-text pairs, we conduct a comparative analysis of our approach against CLIPSelf [46] in utilizing these pairs. As shown in Tab. 11, CLIPSelf employs the substitution of random image crops with pseudo region-text pairs, resulting in an enhanced recognition for foreground objects while concurrently observing a reduction in the accuracy of background identification. In contrast, our proposed DenseVLM achieves a notable improvement in the recognition accuracy of foreground objects while also improving the identification of background stuff.
@@
ViT-L/14 DenseVLM 76.2 92.9 73.3 87.3 47.4 79.1 54.0 84.1 54.2 79.9 57.8 85.9
ViT-L/14 DenseVLM* 75.2 91.8 73.3 87.1 45.5 78.1 54.5 85.0 55.6 82.1 58.1 86.4
Table 12: Ablation study on various backbones. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks. methodname* indicates the model initialized by EVA-CLIP [41].
@@ Appendix D Visualizations
Confusion matrix. We compare the region classification results of our method against previous approaches through a confusion matrix visualization for panoptic masks (both thing and stuff categories) on the COCO Panoptic dataset. These confusion matrixs offer a systematic overview of region classification performance, illustrating the incidence of accurate and erroneous classifications, particularly facilitating a precise assessment of models’ accuracy in differentiating between thing and stuff categories. As shown in Fig. 6, Fig. 7, and Fig. 8, prior methods, including EVA-CLIP [41], RegionCLIP [57], and CLIPSelf [46], often misclassify background regions as co-occurring foreground classes, such as incorrectly identifying snow as skis or sky as kite. In contrast, as demonstrated in Fig. 9, our DenseVLM achieves higher accuracy in recognizing each category, with a notable improvement in the precision of background object identification.
Image grid patches classification. We visualize the classification results of image grid patches using the powerful ViT-L/14 model from CLIPSelf [46]. As shown in Fig. 10, the model demonstrates a stronger focus on foreground object recognition, despite substantial portions of background patch regions being misclassified as foreground objects. The lightweight VLMs are prone to learning these errors. Furthermore, regions with incorrect classifications typically have low confidence scores, making it essential to filter out these regions.

### v2

@@ 3.2 DenseVLM framework
Figure 4: Mask accuracy comparison across categories in COCO dataset. Our method achieves notable improvements, especially in addressing foreground bias. The foreground categories are shown in black, and the background categories are highlighted in red.
Decoupled region alignment to train VLMs without foreground bias. Upon the established region-category relations, we perform alignment between the region and text representations to train the U-VLM. A straightforward approach would be to directly align region features and text embeddings of their corresponding categories while maximizing the separation of unrelated categories. However, since the U-VLM directly inherits from the P-VLM, our experimental results in Sec. 4.3 reveal that this method exacerbates foreground bias, leading to improved foreground detection but limited gains in background recognition.
To mitigate this issue, we propose a decoupled alignment strategy that separates the alignment process for foreground and background regions. Specifically, we denote the region features of the training U-VLM as $\mathcal{F}_{U}$, using the same partitioning grid as the P-VLM. The text embeddings are represented as $\mathcal{T}_{U}$. By leveraging the region-category relations $(k,c^{k})$ retrieved by the P-VLM, we establish a one-to-one mapping for the U-VLM to associate the region features and their corresponding category embeddings. To distinguish the semantic regions of foreground and background, we decouple these region-category relations into two distinct groups, following the predefined two category sets: foreground objects Thing ($\mathcal{U}$) and background regions Stuff ($\mathcal{V}$). By selectively contrasting against text embeddings related to irrelevant categories, we guide the model to focus more on the relevant background regions, reducing the impact of irrelevant foreground categories. This selective contrast helps the model capture the distinctive characteristics of background regions, leading to a more accurate separation between foreground and background. The alignment process can be effectively achieved by maximizing the cosine similarity for the region features and the text embeddings. According to Eqn. (2), when $c^{k}\in\mathcal{V}$, the probability $q^{k}$ for a specific region is calculated as:
$q^{k}=\frac{\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}\cup\mathcal{V}}\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{j})/\tau)}.$ (3)
$\tilde{q}^{k}=\begin{cases}\frac{\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}}\exp(\cos(\mathcal{F}^{k}_{U},\mathcal{T}_{U}^{j})/\tau)}&\text{if}~{}c\in\mathcal{U},\\ 0&\text{otherwise}.\end{cases}$ (4)
The overall loss for each image is computed as $\mathcal{L}=\frac{1}{m^{\prime}\times n^{\prime}}\sum^{m^{\prime}\times n^{\prime}}_{k}\mathcal{L}^{k}$, where the sum is taken over all regions, excluding those removed by the region denoising scheme.
4 Experiments
VLMs Region Alignment GPU Memory Time Overhead Boxes Masks-T Masks-S
ViT-B/16 & ViT-B/16 Images Features KD 37G 25min 69.1 88.2 66.7 83.4 41.7 75.2
ViT-L/14 & ViT-B/16 Images Features KD 39G 37min 24.2 52.4 23.4 51.1 10.1 39.1
ViT-L/14 & ViT-B/16 Images Logics KD 39G 55min 72.2 89.8 68.8 83.8 42.6 75.2
ViT-L/14 & ViT-B/16 Features DenseVLM 39G 23min 73.4 90.5 71.0 84.8 45.6 77.8
[…]
4.1 Benchmarks
Experiment settings. To verify the effectiveness of the proposed DenseVLM, we perform experiments on dense prediction tasks using the COCO panoptic [31] val2017 split and ADE20K panoptic [61] val split. Following CLIPSelf [48], we evaluate box classification using pooled box features (Boxes) and mask classification with pooled mask features, distinguishing foreground objects (Masks-T) from background content (Masks-S). This process follows a similar procedure as illustrated in Fig. 1, where we use ground truth annotations to extract local features and evaluate the local classification. Results are reported in Top-1 and Top-5 mean accuracy across all experiments.
Implementation details. We employ the ViT-L/14 model from CLIPSelf [48] as the powerful P-VLM and ViT-B/16 from EVA-CLIP [45] as the training U-VLM. To enhance computational efficiency, the powerful VLM are kept frozen, and only the image encoder of the U-VLM is trained with pre-extracted text embeddings. Considering the practical application of the downstream tasks and the trade-off between performance and efficiency, we resize the input images to a uniform resolution of $512\times 512$ pixels. The models are trained for 6 epochs using the AdamW [36] optimizer with a weight decay of 0.1. The initial learning rate is set to 1×$\times 10^{-5}$ with a cosine decay [35] schedule.
4.2 Comparison with other VLMs
Quantitative evaluation. We perform a comprehensive quantitative evaluation of dense representations across multiple VLMs on the COCO Panoptic [31] and ADE20K Panoptic [61] datasets. As shown in Tab. 1, while previous methods [3, 10, 58, 45] achieve strong zero-shot image classification performance, their performance in region recognition is notably suboptimal. For instance, EVA-CLIP [45] achieves only a Top-1 box classification accuracy of 44.3% on COCO and 33.0% on ADE20K. Although RegionCLIP [60], which is trained on region-text pairs, shows improved performance on COCO, but fails to generalize to datasets such as ADE20K. Moreover, FineCLIP [20] and CLIPSelf [48], which incorporate self-distillation, achieve higher Mask-T classification accuracy but perform poorly in Mask-S classification. In contrast, our method, DenseVLM, notably outperforms these models, achieving a 4.3% improvement in Top-1 Mask-T accuracy and a 3.9% improvement in Top-1 Mask-S accuracy on COCO, highlighting its effectiveness in dense prediction tasks.
Fig. 4 shows a comparison of mask accuracy across different categories, further demonstrating the effectiveness of DenseVLM in region-text alignment. Notably, our method notably improves accuracy in background categories such as ‘sky’ and ‘wall’, which can mitigate foreground bias.
[…]
Qualitative results. We visualize attention maps using cosine similarity for text-described object categories. As shown in Fig. 5, DenseVLM achieves more accurate and complete object localization than EVA-CLIP and CLIPSelf. Moreover, it better preserves semantic separation, reducing interference from other objects.
4.3 Ablation study
Framework design exploration. We conduct ablation studies to explore key design choices in DenseVLM, including target VLMs for aligning representations, region cropping, and optimization strategies with a focus on GPU memory usage and time overhead. Using a self-distillation strategy [48] as the baseline, we observe a significant performance drop when replacing the target model with ViT-L/14 CLIP from CLIPSelf, due to disruption in visual-language alignment. Replacing feature distillation with logit distillation improves performance but decreases training efficiency due to repeated feature extraction, while introducing foreground bias issues. In contrast, DenseVLM achieves superior performance with a more efficient feature-cropping strategy and decoupled alignment framework, reducing both training time and GPU memory usage.
P-Thing P-Stuff
Thing Stuff Thing Stuff Boxes Masks-T Masks-S
74.3 70.9 42.6
74.2 70.8 42.0
74.1 70.9 41.3
73.4 71.0 45.6
Ablation study on decoupled alignment. Tab. 3 analyzes the effect of the decoupled alignment strategy on DenseVLM. The retrieved category $c^{k}$ can be categorized into Thing and Stuff, referred to as ‘P-Thing’ and ‘P-Stuff’, respectively. DenseVLM leverages $c^{k}$ for regions to selectively contrast against irrelevant Thing and Stuff categories. When regions contrast against all categories, the model can better identify the foreground regions but struggles with distinguishing the background (1st row). In the fully decoupled setting, where each region contrasts against its own set of categories, performance further degrades (2nd row). The most pronounced foreground bias occurs when P-Thing contrasts against Stuff but P-Stuff does not contrast against Thing (3rd row). In DenseVLM, we adopt a balanced strategy where P-Stuff contrasts against both Thing and Stuff, while P-Thing does not contrast against Stuff, resulting in improved overall performance by mitigating bias (4th row).
[…]
@@ 4.3 Ablation study
Ablation study on different category sets. Tab. 4 presents the effects of different category sets used for retrieval and alignment performance. When using the category set from dataset annotations, we observe that as the number of categories increases, performance consistently improves. Adding only the background category helps the model distinguish foreground differences, as seen in rows 1 and 2, and incorporating more foreground categories leads to performance gains, as seen in rows 3 and 4. These improvements can be attributed to the enhanced representational capacity afforded by a broader category set.
To address the limitations of category set annotations, we generate category sets using the generative model NXTP [56] based on images and use DeepSeek-R1 [14] to classify foreground and background categories. As shown in Tab. 4, our model still achieves strong performance.
Boxes Masks-T Masks-S
Categories Top1 Top5 Top1 Top5 Top1 Top5
The category set is from the dataset annotation.
133 (80) 71.1 88.5 68.7 83.0 44.7 75.2
171 (80) 72.3 89.8 69.4 85.8 44.2 76.0
273 (160) 72.3 89.9 70.1 84.4 44.9 76.4
316 (204) 73.4 90.5 71.0 84.8 45.6 77.8
The category set is generated by NXTP based on the images.
210 (133) 72.6 90.2 70.4 84.6 41.5 75.7
794 (484) 72.5 90.4 70.3 84.8 44.1 76.2
Ablation study on different P-VLMs. Tab. 5 shows that using ViT-B/16 as the retrieval model, our method performs well in mask classification, outperforming the P-VLM (ViT-B/16 with CLIPSelf). This improvement effectively mitigates foreground bias by enhancing the distinction between foreground and background regions. Using ViT-L/14 further improves performance across all metrics.
@@ 4.4 Zero-shot comparisons at different data scales
Boxes Masks-T Masks-S
P-VLM Top1 Top5 Top1 Top5 Top1 Top5
ViT-B/16 70.0 88.6 68.0 83.0 43.0 75.3
ViT-L/14 73.4 90.5 71.0 84.8 45.6 77.8
Table 5: Ablation study on different P-VLMs utilized for retrieving categories for unlabeled regions.
4.5 Extending to different backbones
4.6 Application to open-vocabulary dense tasks
@@
ViT-B/16* CLIPSelf 67.6 64.4 44.5
ViT-B/16* DenseVLM 71.9 70.0 47.8
ViT-B/16 EVA-CLIP 44.3 44.7 26.2
ViT-B/16 CLIPSelf 69.1 66.7 41.7
ViT-B/16 DenseVLM 73.4 71.0 45.6
@@
Table 6: Results and comparisons of different backbones. * indicates the model initialized by OpenCLIP [3].
Method A-150 A-847 PC-459
PACL [40] 31.4 - -
OVSeg [29] 24.8 7.1 11.0
MAFT [18] 29.1 10.1 12.6
SED [51] 31.6 11.4 18.6
SCAN [33] 30.8 10.8 13.2
CAT-Seg+CLIPSelf [48] 29.7 10.1 -
CAT-Seg+FineCLIP [20] 32.4 12.2 -
SAN [53] 27.4 10.0 13.0
SAN+DenseVLM 29.5+2.1 10.4+0.4 15.6+2.6
CAT-Seg [4] 31.4 11.7 18.4
CAT-Seg+DenseVLM 34.1+2.7 12.2+0.5 18.7+0.3
@@ 5 Conclusions
  • In Sec. C, we offer supplementary experiments on input image sizes, region proposals, backbones, category sets and threshold of region denoising for training VLMs.
@@ Appendix A Limitations and broader impact
Limitations: Our aim is to develop a region-language alignment model that effectively integrates local visual and semantic features, thereby improving open-vocabulary dense prediction performance. Compared to previous pre-trained Vision-Language Models (VLMs) [3, 45, 60, 48], our proposed DenseVLM achieves superior results and significantly improves downstream task performance. We believe DenseVLM has even greater potential. 1) Scalability. DenseVLM is designed within an efficient, unsupervised region-language alignment framework, making it adaptable to various datasets. However, computational resource limitations have restricted our ability to scale to larger datasets. 2) Model capacity. We employ the ViT-L/14 model from CLIPSelf [48] as a powerful Pre-trained VLM (P-VLM). Utilizing more robust VLMs can yield better performance, and transferring their rich semantic knowledge to training models is a promising direction. 3) Fine-grained Semantic. We categorize objects into broad thing and stuff classes. Fine-grained semantic segmentation and decoupled alignment would enhance the model’s ability to distinguish between similar categories. We plan to explore these avenues in our future research.
@@ Appendix C Additional experiments
Ablation study on input image sizes. To evaluate the effect of input image size on DenseVLM, we conduct experiments with distinct resolutions: 224, 320, 512, 768 and 1024 pixels, for both training and inference. As shown in Tab. 10, model’s performance on the region classification task improves as image resolution increases from 224 to 1024 pixels. This enhancement can be attributed to the greater detail captured at higher resolutions. However, this improvement comes a significant increase in GPU memory usage. Considering the trade-off between computational resources and model performance, we resize the images to 512$\times$512 pixels to achieve an optimal balance.
Boxes Masks-T Masks-S
$\theta$ Top1 Top5 Top1 Top5 Top1 Top5
0.0 72.1 89.6 68.2 84.3 43.6 76.1
0.1 72.7 90.2 69.1 84.3 44.6 77.2
0.2 73.1 90.4 69.7 84.6 45.1 77.7
0.3 73.4 90.5 71.0 84.8 45.6 77.8
0.4 73.2 90.2 70.2 84.3 45.2 77.5
0.5 73.1 90.0 70.0 84.3 45.0 77.1
0.6 73.1 89.9 69.6 84.0 44.6 76.3
Table 12: Ablation study on threshold of $\theta$ in region denoising.
Ablation study on using region proposals. Following RegionCLIP [60] for fine-tuning VLMs with pseudo-labelled region-text pairs, we compare our approach to CLIPSelf [48] in utilizing these pairs. As shown in Tab. 11, CLIPSelf substitutes random image crops with pseudo region-text pairs, resulting in an enhanced recognition for foreground objects while concurrently observing a reduction in the accuracy of background identification. In contrast, our proposed DenseVLM achieves a notable improvement in the recognition accuracy of foreground objects while also improving the identification of background stuff.
Ablation study on the threshold $\theta$. We perform an ablation experiment to assess the impact of varying threshold $\theta$ values of region denoising. As shown in Tab. 12, the model performs the worst when $\theta=0$. When $\theta$ is set lower, the Top-5 accuracy increases, but results in suboptimal performance. This may be due to low-confidence categories causing alignment confusion for the model. Conversely, setting $\theta$ too high filters out too many local images, decreasing performance. By defaul, we select $\theta=0.3$ for DenseVLM.
@@
ViT-L/14 DenseVLM* 76.2 92.9 73.3 87.3 47.4 79.1 54.0 84.1 54.2 79.9 57.8 85.9
ViT-L/14 DenseVLM 75.2 91.8 73.3 87.1 45.5 78.1 54.5 85.0 55.6 82.1 58.1 86.4
Table 13: Ablation study on various backbones. We report the Top1 and Top5 mean accuracy on classifying boxes and panoptic masks (thing and stuff) on COCO panoptic and ADE20K panoptic benchmarks. * indicates the model initialized by OpenCLIP [3].
@@ Appendix D Visualizations
Confusion matrix. We compare the region classification results of our method against previous approaches through a confusion matrix visualization for panoptic masks (both thing and stuff categories) on the COCO Panoptic dataset. These confusion matrixes offer a systematic overview of region classification performance, illustrating the incidence of accurate and erroneous classifications, particularly facilitating a precise assessment of models’ accuracy in differentiating between thing and stuff categories. As shown in Fig. 7, Fig. 8, and Fig. 9, prior methods, including EVA-CLIP [45], RegionCLIP [60], and CLIPSelf [48], often misclassify background regions as co-occurring foreground classes, such as incorrectly identifying snow as skis or sky as kite. In contrast, as demonstrated in Fig. 10, our DenseVLM achieves higher accuracy in recognizing each category, with a notable improvement in the precision of background object identification.
Image grid patches classification. We visualize the classification results of image grid patches using the powerful ViT-L/14 model from CLIPSelf [48]. As shown in Fig. 11, the model focuses heavily on foreground object recognition, but significant portions of background patches are misclassified as foreground objects. The training VLMs are prone to learning these errors. Furthermore, regions with incorrect classifications often have low confidence scores, highlighting the importance of filtering them out.